OpenStudy (anonymous):

what is the two equations using (1,6) and (-3,-6).. this is finding the lines of a equations.. formula is y-y=m(x-x), but i dont understand how to get the equations, please explain..

OpenStudy (compassionate):

Okay, first you want to find the slope Slope: m=(y2-y1)/(x2-x1) Then you just plug in the points.

OpenStudy (anonymous):

i know what the slope is, i just dont know how to get the equations...

OpenStudy (compassionate):

So we know that (x, y) is the form they're plugged in, right? We know the two points are: (1,6) and (-3,-6) x = 1 x2 = -3 And y = 6 y2 = -6 Now use the equation: y - y2 = m(x - x2) You should have: 6 - (6) = m(1 - [-3]) All you need is the slope and then add like terms.

OpenStudy (anonymous):

your not understanding what i am asking for..